# DiffArtist: Towards Structure and Appearance Controllable Image Stylization

### v3

@@ 3.4. Structure and Appearance as Delegate Process
We leverage two additional diffusion processes with shared information for dual controllability, which we call delegate branches. We initialize the structure and main branch from the inverted noise, while appearance delegation starts from a sampled noise $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. With these delegations, we are able to control the stylization over the entire diffusion process. The controlled main stylization branch can be denoted as:
@@
(3) $\epsilon^{m}_{\theta}(\mathbf{x}_{t},t;y,\kappa^{s},\psi^{a})=\mathcal{G}_{t}\left(\kappa^{s}\circ\kappa^{m},\psi^{a}\diamond\psi^{m}\right),\quad t\in[1,T]$
@@
where the superscripts $s$, $a$, and $m$ denote the factorization extracted from the structure, appearance delegation, or main branch, respectively. The $\circ$ and $\diamond$ are two non-commutative control operators.
@@ 3.5. Structure and Appearance Representations in Denoising U-Net
In this section we formulate the $\kappa$ and $\psi$ in U-Net-based denoiser for enabling structure and appearance disentangled style control.
@@
Pyramidal Structure Representation. A multi-scale, and semantic-rich image representation is essential for structure stylization at different level of abstractness. To achieve this, we find the hidden features in the ResBlock of denoising U-Net encode appearance-invariant image semantics across varying $t$ (see Fig. 3). Formally, we denote the hidden feature of ResBlock as $h_{i}(\mathbf{x}_{t})$, where $i\in\{1,2,\ldots,N_{res}\}$ indexes the ResBlocks in the decoder with varying spatial resolution up to $N_{res}$. Stacking the hidden feature from all layers forms a feature pyramid as the structure representation of $\mathbf{x}_{0}$ at $t$:
(4) $\kappa_{t}$ :=$:=\{h_{i}(\mathbf{x_{t}})\}_{i\in S_{res}},$
[…]
and $\text{and }S_{res}$$\subseteq\{1,2,\ldots,N_{res}\}.$
Technically, the representation above captures the semantics of $x_{0}$ at multiple spatial resolutions and (almost) full signal-to-noise ratio (SNR), making it ideal for structural stylization that requires semantic awareness. This is fundamentally different with latent-based, solidary representation (e.g., $\mathbf{x}_{\tau}$) and pixel-level representation (e.g., ControlNet, IP-Adapter), which are uni-resolution, at single SNR level, and semantically-poor. Consequently, as evidenced in our experiment, these methods fail to produce meaningful structure styles and are prone to deterioration of original semantics, both in terms of structure and appearance. We implement the content control operator $\circ$ as injection (i.e., $a\circ b=a$).
@@
Semantic-aware Appearance Representation. Akin to structure, the appearance-related style also must also be harmoniously aligned with the image semantics. However, until now, the appearance delegation denoise from a random noise and hence has no interaction with the source image. To solve this, we propose Structure-to-Appearance injection (S2A) to propagate the high-level semantics into the generation of appearance to avoid spatial misalignment. Specifically, we inject self-attention value $V^{s}$ from early layers of $\epsilon^{s}_{\theta}$ to $\epsilon^{a}_{\theta}$.
We represent the appearance of the target style as self-attention maps from all layers in the $\epsilon^{s}_{\theta}$. Let $N_{attn}$ denote the total number of attention blocks within the U-Net decoder, $S_{s2a}\subseteq\{1,2,\ldots,N_{attn}\}$ be the selected blocks for s2a injection. The style representation is:
(5) $\sigma^{a}_{t}:=\{A^{s}_{i}\}_{i\in[1,2,\cdots,N_{attn}]}$
[…]
@@
Figure 3. ResBlock feature map visualization. We apply t-sne to visualize the feature map of different feature maps in U-Net. The hidden features $h(\mathbf{x})$ better preserves the semantics than the ResNet feature $f(\mathbf{x})$ throughout all $T$.
@@
Figure 4. Qualitative comparison with existing methods. We compare our work with representative text-driven image manipulation method in (a), and image-based stylization methods in (b). Stylized images generated by DiffArtist produce high-fidelity structural and appearance-level style with semantic integrity. We suggest readers for more visualizations in Appendix. A.
@@
(6) $a\diamond b=\operatorname{Var}(b)(\frac{a-\operatorname{Mean}(a)}{\operatorname{Var}(a)})+\operatorname{Var}(b).$
The entire control process is visualized in Fig. 2. Adjusting the control layer $S_{res}$ and style strength in $\epsilon_{t}^{s}$ enables disentangled control for structure and appearance, respectively.
3.6. MLLMs as Stylization Evaluator
Recent research demonstrates a strong semantic-level multimodal understanding of MLLMs (Zhang et al., 2024; Kojima et al., 2022; Zhang et al., 2023c). Inspired by these findings, we leverage state-of-the-art MLLMs as strong zero-shot artistic evaluators. Due to the inherent subjectivity in artistic evaluation, we argue that obtaining an absolute score for structure or appearance fidelity can be challenging. Consequently, we evaluate them in a relative way, where the results from all compared methods are visible to MLLM for ranking. Besides, we aim to measure the fidelity, instead of the more subjective visual-appeal [^1] Note that, visual-appeal is also not equivalent to the concept of aesthetics. (Gombrich and Gombrich, 1995).
@@ 4.1. Experiment Setup
Implementation details. The pretrained Stable Diffusion 2.1  [^2] https://huggingface.co/stabilityai/stable-diffusion-2-1 is used in our main experiment. We perform DDIM sampling with $T=50$ steps. During the inversion, the intermediate noise predictions are recorded, which are used to overwrite the input of denoiser of the content delegation at every steps. Further implementation details can be found in Appendix B.
We default $S_{\text{res}}$ to be the first four ResNet layers ([1, 2, 3, 4]), and $S_{\text{s2a}}$ as the first two attention layer features ([1, 2]), unless otherwise specified. The classifier-free guidance (CFG) scale is set as $7.5$. Note that the default parameter corresponds to only moderate structural variations, used to set a fair comparison with other works to avoid per-image tuning. However, it should be noted that this paper focuses on improved controllability to allow for customization, allowing user to personalize according to their taste. Experiments are conducted with a single RTX 4090-D GPU. The inversion process takes approximately 2 seconds, and the subsequent stylization process takes around 8 seconds.
Compared methods. We compare our method against existing text-driven stylization and manipulation methods: DDIM Inversion (Song et al., 2020), CLIPStyler (optimization-based) (Kwon and Ye, 2022), DiffStyler (Huang et al., 2024), Plug-and-Play (PnP), Prompt2Prompt (with null text inversion) (P2P)  (Mokady et al., 2023; Hertz et al., 2022), ControlNet (Zhang et al., 2023a), and InstructPix2Pix (Brooks et al., 2023). Additionally we consider a baseline named SD, which generate image with the Stable Diffusion according to $y$.
We also indirectly compare our method with reference-based stylization methods, including AdaIN (Huang and Belongie, 2017), ArtFlow (An et al., 2021), StyTr2 (Deng et al., 2022), StyleID (Chung et al., 2024), StyleAligned (Hertz et al., 2024) (with ControlNet) and InstantStyle (Wang et al., 2024a) (with ControlNet). Images generated by SD are used as reference.
Conventional Metrics: LPIPS (Zhang et al., 2018) measures the content preservation by calculating the perceptual distance from source and stylized image. For style fidelity, we leverage CLIP Score (Radford et al., 2021) and Pick Score (Kirstain et al., 2023), both quantify the alignment between the stylized image $\hat{I}$ and prompt $y$. We also include a human study crowd-sourced from $n_{1}=200$ users and report the average preference rate.
MLLM-based Metrics: We rank the stylization fidelity in ascending order, and normalize the rankings across all test samples. Therefore, a score closer to 1 indicate stronger fidelity. We use Gemini-v2.0-flash is used due to its strong multimodal task-solving capability and empirically superior alignment with human. The full prompt templates can be found in Appendix C.
@@ 4.2. Comparisons
Qualitative Comparisons. We first provide a comprehensive comparison against previous text and reference-based stylization methods, visualized in Fig. 4-(a, b). (a): When compared with text-driven methods, DiffArtist is the best at following the style prompt while maintaining content integrity. Our method enables harmonious structural variations, such as pixelation, without compromising intricate details like facial identity and hair. In contrast, the compared methods (e.g., DiffStyler, Plug-and-Play) may introduce undesired modifications that violates semantics or produce misaligned styles (e.g., InstructP2P, ControlNet). (b): When broadly compared with reference-based methods, our method still stands out for its capability to curate both appearance and structural variations. To fully demonstrate the results, we suggest readers for additional visualizations in Appendix A.
Quantitative Comparison. For our quantitative evaluation, we first sample 50 art styles from WikiART, with both abstract (e.g., “Cubism”) and realistic styles (e.g, “High-renaissance”), which are diversified by GPT-4o in terms of description. This sets a wide spectrum of styles to align with real-world applications. The content comprises 50 images from MSCOCO (Lin et al., 2014) and 50 photorealistic images generated by another model (Ideogram, 2024). For each content image we randomly draw $10$ style prompts from all possible styles, resulting in a total of $1,000$ combinations for comparison. The comparative results are presented in Tab. 1.
For conventional metrics, DiffArtist achieves an LPIPS of 0.52, a CLIP Score of 25.91, and a PickScore of 20.51, outperforming most of the compared methods. However, these metrics discard either the content or style reference, and they do not explain stylization quality in terms of structure and appearance. For example, the baseline SD has be best CLIP Score and PickScore but it completely discard the content information in the original image. We include these scores solely for reference.
When evaluated with MLLM-based metrics, DiffArtist attains the highest average score. Specifically, our method achieves the second-highest structure score of 0.61, demonstrating DiffArtist’s effectiveness in generating structural styles. The only stronger baseline is P2P, which is designed for image editing, with relatively weak stylization capability. Besides, our method achieves the best appearance score, showcasing the advanced capability of DiffArtist in capturing high-fidelity appearance details embedded in style prompt. In human evaluations, our method is preferred 67.8% of users, further validating its superiority.
@@ 5. Analysis and Discussion
This section provided extended comparisons and discussions. In Sec. 5.1 we provide an in-depth comparison of different style control methods, both qualitative and quantitatively. Sec. 5.2 validate the effectiveness of MLLMs as style evaluator. Sec. 5.3 we provide ablations on delegation and S2A injection. We conclude this section with a discussion in Sec. 5.4.
@@
Figure 5. Comparisons between representative style control methods. DiffArtist offers significantly stronger controllability than other stylization methods. (Left): DiffArtist allows smooth and artistically meaningful structural stylization strength control, while maintaining appearance style strength. (Right): DiffArtist allows fine-grained control of appearance-related style strength while preserving structural and semantic integrity.
5.1. DiffArtist has Fine-grained Controllability
@@
This section aims to validate the unprecedented dual controllability of DiffArtist. Specifically, we provide fine-grained comparisons of the following three groups of representative methods, configured with different control levels: (a) Semantic pyramid: include DiffArtist and DiffArtist implemented with Plug-and-Play structure representation, $f(\mathbf{x})$ (Ours + PnP), and Ctrl-X; (b) Pixel-level map: include InstantStyle (Cui et al., 2023), which is based on ControlNet (Zhang et al., 2023a) to use pixel-level maps as additional conditioning; (c) Noise inversion, which corresponds to the DDIM baseline. For content control in group (a), we set 5 levels, using the following control layers: ($\varnothing$, [1],[1-4],[1-6],[1-8]); for group (b), we evenly sample their respective control strength hyperparameters; and we sample $\tau=[0,5,10,15,20]$ for group (c). Style strength is controlled by sampling CFG weights in [2.5,5,7.5,10] for all groups except for Ctrl-X, which is achieved by adjusting its appearance schedule parameter.
Qualitative Comparison. As visualized in Fig. 5, we observe the strength of structure and appearance style are the most adjustable in DiffArtist, with harmonious, consistent, and disentangled interpolations across a sequence of control levels. Specifically, the compared methods (Ours+PnP) and group (b) primarily produce style at the appearance level, failing to capture the essential fragmented shape of Cubism at varying levels of abstraction. When tuned with different appearance strength, these methods are also incapable of producing artistically meaningful interpolations. While DiffArtist produce gradually bolder strokes, the compared methods only gives more saturated results. This experiment also demonstrate the effectiveness of content representation: the pixel-map based representation can severely violate the semantic integrity, like the example of InstantStyle where the face is barely recognizable. The inversion-based representation, on the other hand, entangles structure and appearance, making it difficult to control one without affecting the other. Further visualization in Appendix. A more clearly demonstrate the strength of DiffArtist.
@@
Quantitative Comparisons. We provide quantitative experiments to further validate the observations from the visualization above. We consider two properties of control: quality and editability. (1) We compare the quality, measured as structure and appearance fidelity scores at different control levels, across different methods. The results are reported in Tab. 2. DiffArtist outperforms others significantly and consistently, demonstrating superior control quality.
We further consider (2) the editability and disentanglement properties of the control within each method. For instance, when increasing structure control, one would expect a gradually stronger structure style, covering a wide range (editability) with minimal influence on the appearance-related style (disentanglement), and vice versa. In other words, during a sequence of control, the structure fidelity should exhibit monotonic growth while the appearance score should resemble a random distribution. To quantify this intuition, we use Spearman’s coefficient of correlation $\rho$ to measure the monotonicity of control, and Kendall’s coefficient of concordance $W$ to quantify the randomness of sequential manipulation. A $\rho$ close to 1 indicates stronger monotonic correlation, while a $W$ close to 0 indicates random ranks. As shown in Tab. 2, the control of DiffArtist is the most editable and disentangled, reaffirming the observation of the visualizations in Fig. 5.
@@
Table 2. Fidelity of structure and appearance control via cross-method comparison. The values corresponds to structure and appearance score (MLLM, $\uparrow$), respectively. Note that the magnitude of scores are only comparable within each column. The best result for each column is in bold.
@@
Table 3. Control editability and disentanglement via inter-method comparison. Higher $\rho$ indicates stronger correlation while lower $W$ means ranks are random. When controlling from one perspective, a high $\rho$ is desired for editability, and a low $W$ for the other aspect is expected for disentangled control. The controls in DiffArtist are the most editable and disentangled.
@@
5.2. MLLMs are Human-Aligned Style Transfer Metrics
We evaluate how each stylization metric is aligned with human perception by calculating the statistical correlation with rankings from human feedback. To achieve this, we first construct a comparison set of 800 stylized images [^3] The images tested here do not overlap with the experiment in Tab. 1, and compare how human and MLLM preferences correlate.
For human feedback, we consider two groups of users. For the non-expert group, we distributed questionnaires on a crowd-sourced platform, where each participant was required to complete randomly sampled ranking tasks. A total of $n_{1}=200$ participants took part in this study. Filters were applied to validate and clean the raw feedback. We also recruited an expert group of $n_{2}=12$ participants (including the author(s)), who have experience in painting or creative design.
We utilize spearman’s $\rho$ to measure the alignment between metrics and human feedback. The averaged (of all content-style pairs) coefficient $\rho$ for both groups is reported in Tab. 4. As the table shows, the MLLM-based metrics are better aligned with human perception, suggesting its effectiveness as an evaluation metric for style transfer.
Table 4. Metrics correlation with human feedback. We report correlation $\rho$ and combined significance $p$. The VLM-metrics align better with human evaluation for both expert and non-expert users.
Corr. (Non-expert) Corr. (Expert) Metrics $\rho$uparrow p-value $\downarrow$ $\rho$uparrow p-value $\downarrow$ S SSIM (Wang et al., 2004) 0.29 0.12 0.25 0.14 MLLM (GPT-4o) 0.44 0.004 0.34 0.20 MLLM (Gemini v2) 0.42 0.02 0.45 0.03 A CLIP Score 0.05 0.73 0.01 0.75 Pick Score (Kirstain et al., 2023) 0.27 0.11 0.25 0.13 MLLM (GPT-4o) 0.25 0.05 0.22 0.06 MLLM (Gemini v2) 0.48 0.04 0.41 0.02

### v4

@@ 3.4. Structure and Appearance as Delegate Diffusion Process
To break this bottleneck, we propose a novel mechanism that stylizes an image with separate diffusion trajectories, as illustrated in Fig. 2. Specifically, we leverage two supplementary diffusion processes with shared information, called delegate branches. We initialize the structure and main branch from the inverted noise $\mathbf{x}_{T}$, while appearance delegation starts from a Gaussian $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. These delegations enable controlling the stylization over the entire diffusion process. The controlled main branch can be denoted as:
@@
(3) $\epsilon^{m}_{\theta}(\mathbf{x}_{t},t;y,\kappa^{s}_{t},\psi^{a}_{t})=\mathcal{G}_{t}\left(\kappa^{s}_{t}\circ\kappa^{m}_{t},\psi^{a}_{t}\diamond\psi^{m}_{t}\right),\quad t\in[0,T]$
@@
where the superscripts $s$, $a$, and $m$ denote the factorization extracted from the structure, appearance delegation, and main branch, respectively. The $\circ$ and $\diamond$ are two non-commutative control operators.
@@ 3.5. Structure and Appearance Representations in Denoising U-Net
Having established the control mechanism in Eq. 3, we now formulate the $\kappa$ and $\psi$ in a U-Net-based denoiser for disentangling structure and appearance control.
@@
Pyramidal Structure Representation $\kappa$. To effectively control structural stylization, we require a representation that captures image semantics at multiple levels of abstraction. We identify the hidden features in the ResBlock of denoising U-Net as the ideal substrate for this purpose, which robustly encode appearance-invariant image semantics across varying $t$ (see Fig. 3). Formally, we denote the hidden feature of a ResBlock as $h_{i}(\mathbf{x}_{t})$, where $i\in\{1,2,\ldots,N_{res}\}$ indexes the ResBlocks up to $N_{res}$, with increasing spatial resolution. Stacking such feature from all layers forms a pyramidal structure representation of $\mathbf{x}_{0}$ at $t$:
(4) $\kappa^{s}_{t}$ $=\{h_{i}(\mathbf{x_{t}})\}_{i\in S_{res}},$
$\text{where }h_{i}$ $\text{ extracted from }\epsilon^{s}_{\theta}(\mathbf{x}_{t},t;\varnothing),$
$\text{and }S_{res}$ $\subseteq\{1,2,\ldots,N_{res}\}.$
Our representation is distinct as it captures multi-scale semantics and provides continuous guidance across the full denoising trajectory ($t\in[0,T]$). This fundamentally differs from methods relying on solitary control points (e.g., $\mathbf{x}_{\tau}$) or single-scale conditions (e.g., ControlNet, IP-Adapter). Such approaches are constrained to a fixed resolution and/or SNR, which architecturally limits their ability to generate complex structural styles. As evidenced in Sec. 5, this constraint often leads to undesirable semantic trade-offs. With this pyramidal representation, we implement the structure control operator $\circ$ as injection (i.e., $a\circ b=a$).
@@
Semantic-aware Appearance Representation $\psi$. We represent the appearance of the target style as self-attention maps extracted from all layers of $\epsilon^{s}_{\theta}$. For the style to be applied harmoniously, its generation must be guided by the image’s semantics. However, until now, we denoise appearance delegation from Gaussian noise and hence has no information-sharing with the source image. To compensate for this, we propose Structure-to-Appearance injection (S2A) that propagates the high-level semantics into appearance generation. Specifically, we inject the self-attention value $V$ from early layers of $\epsilon^{s}_{\theta}$ to $\epsilon^{a}_{\theta}$. Let $N_{attn}$ denote the total number of attention blocks within the U-Net decoder, $S_{s2a}\subseteq\{1,2,\ldots,N_{attn}\}$ be the selected blocks for S2A injection. The appearance representation at $t$ is:
(5) $\psi^{a}_{t}=\{A^{a}_{i}\}_{i=1}^{N_{attn}},\quad\text{where }$ $A^{a}_{i}\text{ is extracted from }\epsilon^{a}_{\theta}\left(\mathbf{x}_{t},t;y,\{V^{s}_{j}\}_{j\in S_{s2a}}\right),$
with $\{V^{s}_{j}\}_{j\in S_{s2a}}\text{ extracted from }\epsilon^{s}_{\theta}(\mathbf{x}_{t},t;\varnothing).$
@@
Figure 3. ResBlock feature map visualization. We apply t-sne to visualize the feature map of different feature maps in U-Net decoder. The hidden features $h(\mathbf{x})$ better preserves the semantics than the ResNet feature $f(\mathbf{x})$ throughout all $T$.
@@
Figure 4. Qualitative comparison with existing methods. We compare our work with representative text-driven image manipulation method in (a), and image-based stylization methods in (b). Stylized images generated by DiffArtist produce high-fidelity structural and appearance-level style with semantic integrity. We suggest readers for more visualizations in Appendix A.
@@
(6) $a\diamond b=\sigma(b)\left(\frac{a-\mu(a)}{\sigma(a)}\right)+\mu(b).$
We visualize the feature interactions in the right part of Fig. 2. Adjusting the control layer in $S_{res}$ and style strength of $\epsilon_{t}^{a}$ enables disentangled control for structure and appearance, respectively.
3.6. Structure and Appearance Evaluation via MLLMs
Recent research demonstrates the powerful semantic-level multimodal understanding of MLLMs (Zhang et al., 2024; Kojima et al., 2022; Zhang et al., 2023c; Li et al., 2025a, b). We leverage state-of-the-art MLLMs as a zero-shot evaluator to assess two key axes of our method: structure preservation and appearance fidelity. Crucially, our goal is to measure fidelity of stylized images, not subjective qualities like visual appeal. [^1] Note that visual-appeal is also not equivalent to aesthetic quality (Gombrich and Gombrich, 1995; Jiang and Chen, 2025).
@@ 4.1. Experiment Setup
Implementation Details. Our experiments are built upon the publicly available Stable Diffusion 2.1 model [^2] https://huggingface.co/stabilityai/stable-diffusion-2-1. We perform DDIM sampling with $T=50$ steps. During the inversion, we record the intermediate noise predictions to overwrite the input of $\epsilon^{s}_{\theta}$ during denoising. Further implementation details are available in Appendix B. Experiments were conducted with a single RTX 4090-D GPU, with an approximate runtime of 2 seconds for inversion and 8 seconds for the final stylization.
Default Parameters. In main experiment, we default $S_{\text{res}}$ to be the first four ResNet layers ([1, 2, 3, 4]), and $S_{\text{s2a}}$ as the first two attention layer features ([1, 2]). The classifier-free guidance (CFG) scale is set as $7.5$. These default control parameters correspond to moderate structural and appearance variations, used to set a fair comparison with existing works to avoid per-image parameter tuning. However, it should be noted that users can adjust these parameters for customization.
Compared Methods. We compare our method against existing text-driven stylization and manipulation methods: DDIM Inversion (Song et al., 2020), CLIPStyler (optimization-based) (Kwon and Ye, 2022), DiffStyler (Huang et al., 2024), Plug-and-Play (PnP), Prompt2Prompt (with null text inversion) (P2P)  (Mokady et al., 2023; Hertz et al., 2022), ControlNet (Zhang et al., 2023a), and InstructPix2Pix (Brooks et al., 2023). Additionally, we consider a baseline named SD, which generates images with Stable Diffusion according to $y$.
We also indirectly compare our method with reference-based stylization methods, including AdaIN (Huang and Belongie, 2017), ArtFlow (An et al., 2021), StyTr2 (Deng et al., 2022), StyleID (Chung et al., 2024), StyleAligned (Hertz et al., 2024) (with ControlNet), and InstantStyle (Wang et al., 2024a) (with ControlNet). Images generated by SD are used as reference.
Conventional Metrics: LPIPS (Zhang et al., 2018) measures the content preservation by calculating the feature distance between the source and stylized image. For style fidelity, we leverage CLIP Score (Radford et al., 2021) and Pick Score (Kirstain et al., 2023), both of which quantify the alignment between the stylized image $\hat{I}$ and prompt $y$. We also include a human study crowd-sourced from $n_{1}=200$ users and report the average preference rate for our method.
MLLM-based Metrics: We prompt the MLLM to rank the fidelity of $k$ stylized images from best (rank 1) to worst (rank $k$). We normalize the integer ranking and average it over the whole evaluation set. Therefore, a score closer to $1$ indicates a stronger fidelity. We use Gemini-v2.0-flash for its strong multimodal capability. The full prompt templates can be found in Appendix C.
@@ 4.2. Comparisons
Qualitative Comparisons. We first provide a comprehensive comparison against previous methods, visualized in Fig. 4-(a, b). (a): Compared with text-driven methods, DiffArtist is the best at following the style prompt while maintaining semantic integrity. Our method enables harmonious structural variations, such as pixelation, without compromising intricate details like facial identity and hair. By contrast, the compared methods may produce misaligned styles (e.g., CLIPStyler, Plug-and-Play) or introduce undesired modifications that violate semantics (e.g., DiffStyler, ControlNet). (b): When broadly compared with reference-based methods, DiffArtist still stands out for its high stylization fidelity from two perspectives. To fully demonstrate the superiority, we highly suggest readers for additional visualizations in Appendix A.
Quantitative Comparison. For our quantitative evaluation, we first sample 50 art styles from WikiART, with both abstract (e.g., “Cubism”) and realistic styles (e.g., “High Renaissance”), which are further diversified by GPT-4o in terms of description. This diversification sets a broad spectrum of styles to align with real-world user inputs. The content comprises 50 images from MSCOCO (Lin et al., 2014) and 50 photorealistic images generated by another model (Ideogram, 2024). For each of the $100$ content images, we randomly draw $10$ style prompts from all possible styles, resulting in a total of $1,000$ unique combinations for comparison. Tab. 1 presents the results.
For conventional metrics, DiffArtist achieves an LPIPS of 0.52, a CLIP Score of 25.91, and a PickScore of 20.51, outperforming most of the compared methods. However, these metrics do not measure stylization quality in structure and appearance. As a simple counter-example, the baseline SD has the highest CLIP Score and PickScore, whereas it is not even performing stylization. We include these scores solely for reference.
When evaluated with MLLM-based metrics, DiffArtist attains the highest average score of 0.64. Specifically, our method achieves the second-highest structure score of 0.61, demonstrating DiffArtist’s effectiveness in generating structural styles. While the editing-focused P2P method scores higher, it does so by sacrificing stylistic strength, evidenced in qualitative comparisons. Besides, our method achieves the best appearance fidelity score, confirming its superior ability to render the appearance details from the text prompt. In human evaluations, our method is preferred by an average of 67.8% of users in pairwise comparison, further validating its superiority.
@@ 5. Analysis and Discussion
This section provide in-depth analysis on the proposed system. In Sec. 5.1, we analyze the controllability of DiffArtist in detail. Sec. 5.2 validates the effectiveness of MLLMs as style evaluators. Sec. 5.3 provides ablations on delegations and the S2A injection. We conclude with a discussion in Sec. 5.4.
@@
Figure 5. DiffArtist offers disentangled and stronger controllability. (Left): DiffArtist enables smooth and artistically meaningful structural stylization at varying degree, without violating appearance style strength. (Right): DiffArtist allows fine-grained control of appearance-related style strength while preserving structural and semantic integrity.
5.1. DiffArtist has Strong Controllability
@@
This subsection validates the unprecedented dual controllability of DiffArtist. To achieve this, we conduct a fine-grained comparative analysis against representative methods, categorizing them by their core control mechanism: (a) Semantic pyramid: include DiffArtist, DiffArtist implemented with Plug-and-Play structure representation, $f(\mathbf{x})$ (Ours + PnP), and Ctrl-X; (b) Pixel-level map: include InstantStyle (Cui et al., 2023), which is based on ControlNet (Zhang et al., 2023a); (c) Noise inversion, which corresponds to the DDIM baseline. For structure control, we define five levels from weakest to strongest. For group (a) we use the following control layers: ($\varnothing$, [1],[1-4],[1-6],[1-8]); for group (b), we evenly sample their respective control strength parameters; and we use $\tau=[0,5,10,15,20]$ for group (c). Appearance strength is controlled by sampling CFG weights in [2.5,5,7.5,10] for all groups except for Ctrl-X, which is achieved by adjusting its appearance schedule parameter.
Qualitative Comparison on Control. As visualized in Fig. 5, DiffArtist demonstrates superior controllability, with harmonious, consistent, and disentangled interpolations across a sequence of control levels. It correctly captures the essential geometric principles of Cubism, whereas competing methods merely apply a superficial texture (e.g., Ours+PnP) or fail to produce meaningful structural variations (e.g., InstantStyle). This nuanced control is further evident when adjusting style strength; DiffArtist provides an artistically meaningful interpretation by producing bolder strokes according to style prompt, while other methods resort to simplistic and often undesirable increases in color saturation. Most critically, our approach preserves semantic integrity. It avoids the catastrophic failures of pixel-map methods like InstantStyle, which can render the face unrecognizable, and also prevents the facial structure corruption seen in inversion-based methods that inherently entangle form with appearance. The extended visualizations in Appendix A with diverse style and content further confirm the superiority of DiffArtist.
@@
Quantitative Comparisons on Control. We provide quantitative experiments to substantiate our visual observations, evaluating control based on two properties: fidelity and editability. We measure the fidelity as structure and appearance MLLM scores at different control levels. The results are reported in Tab. 2. DiffArtist outperforms others significantly and consistently, demonstrating superior control fidelity.
The editability defines the quality of the manipulation itself, which we assess via three criteria: range, monotonicity, and disentanglement. An ideal control of stylization should cover a wide range of stylistic effects (a large spread in MLLM scores), exhibit predictable monotonicity (Spearman’s $\rho\approx 1$ for the target attribute), and maintain disentanglement from other attributes. We measure the disentanglement using Kendall’s $W$ on the off-target scores, where a stable, unaffected score sequence yields $W\approx 0$. For instance, when controlling structural strength, $W$ for the appearance score is desired to be near 0. As shown in Tab. 2, the control of DiffArtist is the most editable: it covers the broadest range of effects, demonstrates the strongest monotonicity, and achieves the best disentanglement, reaffirming the superior control visualized in Fig. 5.
@@
Table 2. Fidelity of structure and appearance control via cross-method comparison. The values correspond to the structure or appearance score (MLLM, $\uparrow$). Note that the magnitudes of scores are only comparable within each column. The best result for each column is in bold.
@@
Table 3. Control editability and disentanglement via inter-method comparison. Higher $\rho$ indicates stronger monotonicity while lower $W$ means ranks are indistinguishable. When controlling from one perspective, a high $\rho$ is desired for editability, and a low $W$ for the other aspect is expected for disentangled control. The controls in DiffArtist are the most editable and disentangled.
@@
5.2. MLLMs are Human-Aligned Stylization Evaluators
We evaluate how each stylization metric aligns with human feedback by calculating the statistical correlation with the rankings from human feedback. To achieve this, we first construct a comparison set of 800 stylized images [^3] The images evaluated here do not overlap with the main experiment in Tab. 1, and compare how human and MLLM preferences correlate.
To gather human feedback, we consider two groups of users. For the non-expert group, we recruited a large-scale group of $n_{1}=200$ participants through a crowdsourcing platform. Each participant performed a series of randomly sampled ranking tasks. To ensure the integrity of the collected data, we implemented attention checks and consistency filters to remove unreliable responses. We also recruited an expert group of $n_{2}=12$ participants with knowledge of fine art.
We measured the alignment between each metric’s rankings and the human-derived preferences using Spearman’s rank correlation ($\rho$). The averaged (of all content-style pairs) $\rho$ for both groups is reported in Tab. 4. As the table shows, the MLLM-based metrics are better aligned with human perception, validating its effectiveness as an evaluation metric for style transfer.
Table 4. Metrics correlation with human feedback. We report correlation $\rho$ and combined significance pp. The MLLM scores show stronger alignment with both expert and non-expert users.
Corr. (Non-expert) Corr. (Expert) Metrics $\rho$ $\uparrow$ p-value $\downarrow$ $\rho$ $\uparrow$ p-value $\downarrow$ S SSIM (Wang et al., 2004) 0.29 0.12 0.25 0.14 MLLM (GPT-4o) 0.44 0.004 0.34 0.20 MLLM (Gemini 2.0) 0.42 0.02 0.45 0.03 A CLIP Score 0.05 0.73 0.01 0.75 Pick Score (Kirstain et al., 2023) 0.27 0.11 0.25 0.13 MLLM (GPT-4o) 0.25 0.05 0.22 0.06 MLLM (Gemini 2.0) 0.48 0.04 0.41 0.02